# Euclid Quick Data Release (Q1)

### v2

@@ 6.2 Considerations on training and inference costs
The results from Sect. 5.6 show that although our predictions have a significant overlap with many of the selections presented, it also highlights that the ROC classifier produces similar overlap performance. Given that this boundary is formed by a simple calculation from the data, rather than requiring a complex model and therefore near-zero overhead to calculate, it may present itself as the preferred method. However, the ROC metric is static and its performance is fixed by its mathematical definition. It must be reiterated that the current pipeline for the diffusion model has had minimum optimisations for the data and the classification task and so can be expected to improve with larger training sets, more advanced network architectures, or the future incorporation of multi-wavelength data as conditioning information. The ROC classifier is also limited to outputting a single number. The generative capabilities of the diffusion model, where it outputs a reconstructed image, provides the potential for a much richer set of scientific applications beyond simple classification such as decomposition as was discussed in Sect. 5.7. Therefore, although it has shown good performance, its limited approach restricts the ROC classifier to only serve as a fundamental baseline for comparison in our study.
In regards to the PSF and Srsic flux classifier, although there are no computational costs to the user, traditional parametric fitting is a serial, per-object process that becomes a bottleneck for survey-scale datasets (Tuccillo et al. 2018). Having parallelisable and scalable alternative methods, even if they feature a large one-time training cost, allow for a strategic investment as we cover more of the extragalactic sky. For AGN identification specifically, the simple parametric models may not be well suited to handle complex or merged morphologies. This could explain the worse AGN-based recall seen for this classifier in Fig. 12.
@@ 7 Conclusion
In this paper, we have used a diffusion-based inpainting model to identify AGN and QSOs using VIS images from the Q1 data (Euclid Quick Release Q1 2025) from the Euclid telescope. Our approach focuses on inpainting the brightest pixels within the centre of a galaxy and employs a novel thresholding approach based on the reconstruction errors to differentiate AGN and QSOs from the rest of the galaxy population. Using only VIS images, our method can generate a large, reliable sample of optically-selected AGN without requiring AGN labels ahead of time. Utilising a standard training and inference pipeline with minimal modifications for astronomical data, our method demonstrates high recall and significant overlap with AGN candidates obtained through more traditional methods, across various wavelengths including optical, near-infrared (NIR), mid-infrared, and X-ray.
@@
  • By demonstrating that VIS images alone can generate reliable and extensive samples of AGN and QSOs, our method showcases the impressive capabilities of Euclid’s imaging. Given that the future of the mission will cover $\mathrm{14\,000\;deg^{2}}$, the ability for AGN identification using only VIS allows for the possibility of some science aims to be met without the immediate need for integrating multiple data sources. This can significantly reduce the time and complexity of processing and analysing the data.
@@
  • Although our model performs well without significant optimisation for our use case, we provide insights into the behaviour of the model given the challenges presented with astronomical imaging. We show how the assumptions of machine-learning literature, particularly regarding fixed noise schedules, can result in non-optimal training on data with a high dynamic range and large variations in signal-to-noise. By highlighting these inefficiencies, we identify a clear direction for future progress in both astronomy and machine-learning domains through the development of adaptive, data-aware training strategies.
@@
In conclusion, our results demonstrate the feasibility and effectiveness of using diffusion-based inpainting for AGN and QSO identification. The ability to use VIS images exclusively for selecting these candidates provides a potential strategic flexibility to the Euclid mission while reducing reliance on complex multispectral data integration. By proving the utility of applying our pipelines to the task of AGN identification, we believe it can be expanded and can significantly contribute to a broader range of astronomical applications, such as identifying gravitational lensing or finding transient objects.
@@ Appendix A Masks
As our model’s aim is not to generate large collections of synthetic data, the variety and scientific validity of the produced galaxies are not put under high scrutiny. Instead, we only check that there are no obvious catastrophic mistakes. In the last two columns, we can see that due to the large mask and fewer pixels to guide the reconstruction, some images produce a visible boundary around the image where the combination of fixed and generated pixels is not as seamless as the other images. Given that the masks used for AGN detection are significantly smaller than those presented in Fig. 17, the fraction of pixels that the inpainting is conditioned on will be much higher, providing tighter constraints on the output compared to these examples. Improvements to the scheduler, as discussed in Sect. 6.1, would also reduce such artifacts.
@@ Appendix C Reconstruction metrics
Figure Fig. 19 show the differences in the rate of change measurement between input images and the inpainting output.
@@ Appendix E Morphology examples
We show output examples from various morphology types to showcase how the diffusion model adapts to different galaxy shapes. Edge-on galaxies, spirals and mergers are shown in Fig. 22, Fig. 23 and Fig. 24, respectively.

### v3

@@ 6.2 Considerations for training and inference costs
The results from Sect. 5.6 show that although our predictions have a significant overlap with many of the selections presented, it also highlights that the ROC classifier produces a similar overlap performance. Given that this boundary is formed by a simple calculation from the data, rather than requiring a complex model and therefore near-zero overhead to calculate, it may present itself as the preferred method. However, the ROC metric is static and its performance is fixed by its mathematical definition. It must be reiterated that the current pipeline for the diffusion model has had minimum optimisations for the data and the classification task and so can be expected to improve with larger training sets, more advanced network architectures, or the future incorporation of multi-wavelength data as conditioning information. The ROC classifier is also limited to outputting a single number. The generative capabilities of the diffusion model, where it outputs a reconstructed image, provides the potential for a much richer set of scientific applications beyond simple classification such as decomposition as was discussed in Sect. 5.7. Therefore, although it has shown good performance, its limited approach restricts the ROC classifier to only serve as a fundamental baseline for comparison in our study.
In regards to the PSF and Srsic flux classifier, although there are no computational costs to the user, traditional parametric fitting is a serial, per-object process that becomes a bottleneck for survey-scale datasets (Tuccillo et al., 2018). Having parallelisable and scalable alternative methods, even if they feature a large one-time training cost, allows us to make a strategic investment as we cover more of the extragalactic sky. For AGN identification specifically, the simple parametric models may not be well suited to handle complex or merged morphologies. This could explain the worse AGN-based recall seen for this classifier in Fig. 12.
@@ 7 Conclusion
In this paper, we have used a diffusion-based inpainting model to identify AGN and QSOs using VIS images from the Q1 data (Euclid Quick Release Q1, 2025) from the Euclid telescope. Our approach focuses on inpainting the brightest pixels within the centre of a galaxy and employs a novel thresholding approach based on the reconstruction errors to differentiate AGN and QSOs from the rest of the galaxy population. Using only VIS images, our method can generate a large, reliable sample of optically selected AGN without requiring AGN labels ahead of time. Utilising a standard training and inference pipeline with minimal modifications for astronomical data, our method demonstrates high recall and significant overlap with AGN candidates obtained through more traditional methods, across various wavelengths including optical, near-infrared (NIR), mid-infrared, and X-ray.
@@
  • By demonstrating that VIS images alone can generate reliable and extensive samples of AGN and QSOs, our method showcases the impressive capabilities of Euclid’s imaging. Given that the future of the mission will cover $\mathrm{14\,000\;deg^{2}}$, the ability for AGN identification using only VIS allows for the possibility of some science aims to be met without the immediate need to integrate multiple data sources. This can significantly reduce the time and complexity of processing and analysing the data.
@@
  • Although our model performs well without significant optimisation for our use case, we provide insights into the behaviour of the model given the challenges presented with astronomical imaging. We show how the assumptions of machine-learning literature, particularly regarding fixed noise schedules, can result in non-optimal training on data with a high dynamic range and large variations in S/N. By highlighting these inefficiencies, we identify a clear direction for future progress in both astronomy and machine-learning domains through the development of adaptive, data-aware training strategies.
@@
In conclusion, our results demonstrate the feasibility and effectiveness of using diffusion-based inpainting for AGN and QSO identification. The ability to use VIS images exclusively for selecting these candidates provides a potential strategic flexibility to the Euclid mission, while reducing reliance on complex multi-spectral data integration. By proving the utility of applying our pipelines to the task of AGN identification, we believe it can be expanded and can significantly contribute to a broader range of astronomical applications, such as identifying gravitational lensing or finding transient objects.
@@ Appendix A Masks
As our model’s aim is not to generate large collections of synthetic data, the variety and scientific validity of the produced galaxies are not put under high scrutiny. Instead, we only check that there are no obvious catastrophic mistakes. In the last two columns, we can see that due to the large mask and fewer pixels to guide the reconstruction, some images produce a visible boundary around the image where the combination of fixed and generated pixels is not as seamless as the other images. Given that the masks used for AGN detection are significantly smaller than those presented in Fig. 17, the fraction of pixels that the inpainting is conditioned on will be much higher, providing tighter constraints on the output compared to these examples. Improvements to the scheduler, as discussed in Sect. 6.1, would also reduce such artefacts.
@@ Appendix C Reconstruction metrics
Figure 19 shows the differences in the ROC measurement between input images and the inpainting output.
@@ Appendix E Morphology examples
We show output examples from various morphology types to showcase how the diffusion model adapts to different galaxy shapes. Edge-on galaxies, spirals and mergers are shown in Figs. 22, 23, LABEL:, and 24, respectively.